OpenStudy (anonymous):

a line passes through the point (2,8) and has a slope of -5. write an equation for this line

OpenStudy (anonymous):

y=mx+b m=slope 8=-5(2)+b solve for b b=18 y=-5x+18

OpenStudy (anonymous):

y-y1 = m (x-x1)

OpenStudy (anonymous):

y-8= -5(x-2)

OpenStudy (anonymous):

y-8= -5x + 10 y+5x-18= 0